OpenStudy (anonymous):

how do you solve this problem -4√(x+9)=20?

OpenStudy (anonymous):

hahahah finding for medal?

OpenStudy (anonymous):

\[-4\sqrt{x+9}=20\]\[(x+9)^{1/2}=20/-4\]\[(x+9)^{1/2}=-5\]\[x+9=\sqrt{-5}\]\[x=\sqrt{-5}-9\]

OpenStudy (anonymous):

\[(-4)\sqrt{x+9}=20\] \[\sqrt{x+9}=20/(-4)\] \[\sqrt{x+9}=-5\] \[x+9=(-5)^{2}\] \[x+9=25\] \[x=25-9\] \[x=16\]